OpenStudy (anonymous):

Two ordered pairs on the hill are given as (6, 3) and (8, 7). Find the slope of the hill.

OpenStudy (jamesj):

Slope = change in y / change in x = (7 -3)/(8-6) = 4/2 = ...

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Now find the equation using the point slope formula?? how do i do tht

OpenStudy (anonymous):

use \[y-y_1=m(x-x_1)\] with \[m=2,x_1=6,y_1=3\]

OpenStudy (anonymous):

thanx so it wud be y-3=2(x-6)

OpenStudy (jamesj):

Yes but you need to rearrange it so you have y = something in x

OpenStudy (anonymous):

how do i do tht?